OpenStudy (anonymous):

TRUE OR FALSE. The constant of variation k for the direct variation of the following equation is ¾. 4y – 3x = 0

OpenStudy (anonymous):

I'm not sure if the constant of variation is the slope, or the standard deviation squared. If it's slope, it is, in fact, 3/4.

OpenStudy (anonymous):

SO THE STATEMENT IS TRUE?

OpenStudy (anonymous):

...Yes, I suppose it is.

OpenStudy (anonymous):

Yes you are right The equation y = mx + b is the slope-intercept form where m is the slope and b is the y-intercept). The slope of this line is the constant of variation. In other words, in the equation y = mx + b, m is the constant of variation. 4y-3x=0 i.e 4y=3x y=3/4